### v1

# Stratified Expert Cloning with Adaptive Selection for User Retention in Large-Scale Recommender Systems

Abstract.
User retention has emerged as a critical challenge in large-scale recommender systems, significantly impacting the long-term success of online platforms. Existing methods often focus on short-term engagement metrics, failing to capture the complex dynamics of user preferences and behaviors over extended periods. While reinforcement learning (RL) approaches have shown promise in optimizing long-term rewards, they face difficulties in credit assignment, sample efficiency, and exploration when applied to the user retention problem. In this work, we propose Stratified Expert Cloning (SEC), a novel imitation learning framework that effectively leverages abundant logged data from high-retention users to learn robust recommendation policies. SEC introduces three key innovations: 1) a multi-level expert stratification strategy that captures the nuances in expert user behaviors at different retention levels; 2) an adaptive expert selection mechanism that dynamically assigns users to the most suitable policy based on their current state and historical retention level; and 3) an action entropy regularization technique that promotes recommendation diversity and mitigates the risk of policy collapse. Through extensive offline experiments and online A/B tests on two major video platforms, Kuaishou and Kuaishou Lite, with hundreds of millions of daily active users, we demonstrate SEC’s significant improvements over state-of-the-art methods in user retention. The results demonstrate significant improvements in user retention, with cumulative lifts of 0.098% and 0.122% in active days on Kuaishou and Kuaishou Lite respectively, additionally bringing several hundred thousand daily active users to each platform.
Retention, Stratified Expert Cloning, Imitation Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
Figure 1. Schematic illustration of our recommendation system based on user state sequences. The system tracks user state transitions $(s_{i},s_{i+1},...)$ and makes content recommendations, i.e., action sequences $(a_{i},a_{i+1},...)$, with the ultimate goal of high user retention reward.
In the era of information abundance, recommender systems have become the cornerstone of online platforms, shaping user experiences and content consumption (Koren et al., 2009; Shi et al., 2017; Hidasi et al., 2015; Quadrana et al., 2018). Among the myriad of challenges faced by modern recommender systems, user retention has emerged as a critical metric that directly impacts long-term growth and sustainability. User retention refers to the ability to keep users consistently engaged and actively interacting with the platform over extended periods. However, optimizing for long-term user retention poses significant difficulties due to the complex dynamics of user preferences and behaviors that evolve over time.
Traditional recommender system approaches, such as collaborative filtering and content-based methods (He and Chua, 2017; Huang et al., 2016; Wang et al., 2019b, a), primarily focus on improving short-term engagement metrics, such as click-through rates or immediate user satisfaction. These approaches often fail to capture the temporal aspects of user interactions and the long-term effects of recommendations on user retention. Consequently, there is a growing need for recommendation strategies that prioritize user retention.
Recent advancements in reinforcement learning (RL) have shown promise in addressing the sequential nature of user interactions and optimizing for long-term rewards (Park and Han, 2013; Cai et al., 2023a; Liu et al., 2024b, a). RL-based methods formulate the recommendation problem as a Markov Decision Process (MDP) to learn policies that maximize cumulative rewards, thereby improving user retention. However, applying RL to optimize long-term user retention in real-world recommender systems remains challenging due to several factors.
Firstly, the delayed impact of recommendations on user retention introduces difficulties in credit assignment and reward attribution. The long-term effects of a recommendation may not be immediately observable, as shown in Figure 1, making it challenging for RL agents to associate specific actions with their corresponding long-term rewards. Secondly, the vast action space of all possible recommendations, coupled with the need for extensive exploration, can lead to sample inefficiency and suboptimal user experiences. Exploring irrelevant or poor-quality recommendations may frustrate users and hinder retention. Lastly, the complex and dynamic nature of user preferences and behaviors pose challenges in defining suitable reward functions that accurately capture the factors influencing user retention.
To address these challenges, we turn to imitation learning, a paradigm that has achieved notable success in large language model (Ouyang et al., 2022; Sun, 2023), robotics (Florence et al., 2022; Jia and Manocha, 2024) and autonomous driving (Samak et al., 2020; Yildirim et al., 2024) by learning from expert demonstrations. The key insight lies in leveraging abundant interaction data from high-retention users (treated as experts) to circumvent RL’s exploration challenges. By leveraging this valuable resource, we can effectively learn recommendation policies that directly optimize for long-term user retention while circumventing the difficulties faced by RL approaches.
Specifically, we propose Stratified Expert Cloning (SEC), a novel imitation learning framework for optimizing long-term user retention in recommender systems. SEC stratifies expert users into multiple levels based on their retention scores and learns specialized behavior cloning policies for each level. This hierarchical approach enables SEC to capture the nuances in expert user behaviors and preferences across different retention levels.
To adaptively select the most suitable expert policy for each user, SEC introduces an adaptive expert selection mechanism that considers the user’s current state and historical retention level. This mechanism dynamically assigns users to the appropriate expert level, ensuring that the recommendations align with the user’s evolving preferences and retention propensity.
Furthermore, SEC incorporates an action entropy regularization term to encourage recommendation diversity and mitigate the risk of overfitting to a narrow set of expert actions. By promoting exploration within the expert policy space, SEC enhances the generalization ability of the learned policies and improves long-term user engagement.
Through extensive offline experiments and large-scale online A/B tests on two major video platforms with hundreds of millions of daily active users, we demonstrate the significant improvements brought by SEC in user retention, compared to state-of-the-art methods. Our SEC framework offers a principled and practically effective approach for optimizing long-term user retention, providing recommender systems with a powerful tool to enhance user experiences and platform sustainability.
The main contributions of this work are summarized as follows:
  • We propose SEC, a novel imitation learning framework for optimizing long-term user retention in recommender systems by learning hierarchical behavior cloning policies from expert users.
  • We introduce an adaptive expert selection mechanism that dynamically assigns users to the most suitable expert level based on their current state and historical interactions.
  • We incorporate an action entropy regularization term to encourage diverse and generalizable recommendation policies.
  • We conduct extensive offline experiments and online A/B tests on two major video platforms, demonstrating the superiority of SEC over state-of-the-art methods in improving user retention.
Figure 2. Overview of the Stratified Expert Cloning (SEC) framework. Left: During training, the framework learns specialized recommendation policies through behavior cloning from users at multiple high-retention expert levels. The state encoder processes user context and state information, while multiple expert policies are trained simultaneously to capture recommendation strategies from different levels of expert users. Right: At inference time, the framework adaptively selects the most appropriate expert policy based on the current user’s state embedding and historical retention level.
2. RELATED WORK
2.1. User Retention in Recommendation Systems
User retention is a crucial objective in recommendation systems, as long-term engagement often generates greater value than short-term interactions. Traditional methods like collaborative filtering and matrix factorization focused primarily on immediate engagement metrics such as click-through rates but struggled to model the temporal dynamics essential for user retention (Koren et al., 2009; Shi et al., 2017). To address this, recent frameworks like DT4Rec (Zhao et al., 2023b) leverage the Decision Transformer to optimize long-term user retention by modeling sequences of interactions. This approach incorporates an auto-discretized reward prompt that preserves partial order relationships between rewards, enhancing retention modeling over time. Other frameworks, such as IURO (Ding et al., 2023), explicitly model interpretable factors influencing retention, using contrastive learning to identify key content (”aha items”) responsible for sustained user engagement. Similarly, Generative Flow Networks (Liu et al., 2024a) have been applied to model retention by treating it as a probabilistic flow over user sessions, addressing the challenges of sparse and delayed retention signals.
In addition to these frameworks, research on recommendation diversity has shown that increasing content variety helps combat recommendation fatigue and improves user retention. Methods like Maximal Marginal Relevance (MMR) and other diversity-driven approaches (Carbonell and Goldstein, 1998; Park and Han, 2013) have demonstrated that diverse recommendations keep user interest high over time. Studies on customer purchase behavior (Kwon et al., 2020) in industries like fashion have empirically validated that diverse product recommendations not only increase purchase rates but also positively influence long-term retention, highlighting the effectiveness of recommendation diversity in both short-term and long-term engagement.
2.2. Imitation Learning
Reinforcement learning (RL) is effective for optimizing long-term rewards but often faces high sample complexity and the risk of suboptimal exploration, especially in recommendation systems where poor actions can harm user experiences. Imitation Learning (IL)  (Ng et al., 2000; Zare et al., 2024) addresses these challenges by training agents to mimic expert behaviors, avoiding the costly trial-and-error process of RL.
Among IL methods, Behavioral Cloning (BC) is particularly notable for its simplicity and efficiency (Pomerleau, 1991; Ross et al., 2011). Treating the problem as a supervised learning task allows BC to learn policies directly from expert demonstrations. This reduces the need for extensive interactions with the environment and lowers exploration risks. Behavior cloning has been effectively used in supervised fine-tuning and preference learning stages for Large Language Models  (Rafailov et al., 2024; Ethayarajh et al., 2024; Ouyang et al., 2022). Consequently, BC is particularly well-suited for applications that interact with users, such as recommender systems.
Advanced IL methods, such as Inverse Reinforcement Learning (IRL) (Ng et al., 2000) and Adversarial Imitation Learning (AIL) (Ho and Ermon, 2016), further extend IL’s capabilities. IRL infers reward functions behind expert behaviors, enabling generalization to unseen scenarios. Reinforcement Learning from Human Feedback (Ouyang et al., 2022; Wulfmeier et al., 2024) is a specialized form of IRL, where human feedback is treated as the expert guidance. AIL, exemplified by Generative Adversarial Imitation Learning (GAIL), aligns agent and expert distributions through adversarial training, excelling in continuous control tasks. IL’s versatility spans domains like large language model (Sun, 2023; Wulfmeier et al., 2024), robotics, autonomous driving, and game AI, with models like DQfD (Hester et al., 2018) and ChauffeurNet (Bansal et al., 2018) leveraging demonstrations to improve learning efficiency and generalization.
In recommendation systems, IL, particularly BC, holds promise for optimizing user retention by mimicking high-retention user behaviors. To enhance BC, we propose Stratified Expert Cloning (SEC), which stratifies expert users based on retention patterns and learns specialized policies for each group. This approach captures diverse user behaviors, improves long-term retention, and retains BC’s strengths in sample efficiency and interpretability.
3. PRELIMINARIES
3.1. Problem Formulation
We consider a large-scale recommender system where a set of users $\mathcal{U}$ interacts with a set of items $\mathcal{I}$ over multiple sessions. The goal is to provide personalized recommendations that maximize long-term user retention, defined as the likelihood of a user remaining active on the platform over an extended period. User retention is often measured by metrics such as the number of active days within a given time window or the time until the next return to the platform.
Let $\tau=\{(s_{1},a_{1}),(s_{2},a_{2}),\ldots,(s_{T},a_{T})\}$ denote a user trajectory, where $s_{t}$ represents the user state at time step $t$ encoding relevant user information, such as demographics, preferences, and recent interactions, while the action $a_{t}$ represents the recommended item or set of items.
Our objective is to learn a recommendation policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that maximizes the expected long-term user retention:
(1) $\max_{\pi}\mathbb{E}_{u\sim\mathcal{U}}[R_{u}|\pi].$
3.2. Challenges in Optimizing Long-term User Retention
Optimizing long-term user retention in real-world recommender systems presents several technical challenges:
  • Delayed Impact and Credit Assignment: The impact of recommendations on user retention is often delayed, making it difficult to attribute the retention outcome to specific actions. For instance, a user may receive a recommendation for a new category, which sparks their interest and leads to increased engagement in the short term. However, the long-term impact of this recommendation on the user’s likelihood to remain active on the platform may not be immediately observable. This temporal gap between actions and rewards complicates the credit assignment problem.
  • Sample Inefficiency and Exploration: The vast action space of all possible recommendations, coupled with the need for extensive exploration, can lead to sample inefficiency and suboptimal user experiences. Exploring irrelevant or poor-quality recommendations may frustrate users and hinder retention. Therefore, it is crucial to balance exploration and exploitation in a sample-efficient manner.
  • Complex User Preferences and Behaviors: User preferences and behaviors are dynamic and evolve over time. A user’s interests may shift from one category to another over time, or their sensitivity to certain types of recommendations may change based on their current life stage or external factors. Capturing these complex dynamics and adapting to user-specific patterns is crucial for effective retention optimization.
Traditional methods, such as collaborative filtering and content-based approaches (He and Chua, 2017; Huang et al., 2016; Wang et al., 2019b, a), primarily focus on short-term engagement metrics and fail to capture the long-term effects of recommendations on user retention. Reinforcement learning techniques (Park and Han, 2013; Cai et al., 2023a; Liu et al., 2024b, a), while promising, face difficulties in credit assignment, sample efficiency, and exploration when applied to the user retention problem.
3.3. Imitation Learning for User Retention
Imitation learning (Zare et al., 2024), specifically behavior cloning, offers a promising approach to address the challenges in optimizing long-term user retention. Behavior cloning is an imitation learning technique that learns a policy to mimic expert actions given states.
Motivation and Advantages: In large-scale recommender systems, we have access to an abundance of interaction data from high-retention users, who can be considered as experts. By leveraging these expert demonstrations, we can learn recommendation policies that directly optimize for long-term user engagement, churn reduction, and user retention. Imitation learning circumvents the difficulties of explicit reward design and costly exploration faced by RL approaches  (Zare et al., 2024). It enables learning directly from successful real-world examples, avoiding the need for trial-and-error in live systems.
Behavior Cloning Approach: Behavior cloning has demonstrated effectiveness in large language model (Ouyang et al., 2022; Rafailov et al., 2024; Ethayarajh et al., 2024), robotics (Florence et al., 2022; Jia and Manocha, 2024) and autonomous driving (Samak et al., 2020; Yildirim et al., 2024). We aim to adapt it to the user retention problem in recommender systems. Our proposed Stratified Expert Cloning (SEC) framework introduces novel components such as multi-level expert stratification, adaptive expert selection, and action entropy regularization. These innovations enhance vanilla behavior cloning to capture the nuances of user retention dynamics and enable personalized, robust policy learning.
The SEC framework builds upon the strengths of imitation learning while addressing key challenges specific to recommender systems. By learning from high-retention experts in a principled manner, SEC provides a practical and effective approach for optimizing long-term user retention. The following sections detail the technical innovations of SEC and demonstrate its significant performance gains over existing methods.
4. THE PROPOSED FRAMEWORK
4.1. Overview of the SEC Framework
The Stratified Expert Cloning (SEC) framework optimizes long-term user retention in recommender systems by learning from high-retention expert users. SEC consists of three main components. First, multi-level expert behavior cloning stratifies expert users into $K$ levels based on retention scores and learns a specialized policy $\pi_{k}$ for each level. Second, action entropy regularization encourages recommendation diversity by incorporating an entropy regularization term in the training objective. Third, an adaptive selection mechanism assigns users to the most appropriate expert level based on their current state and historical interactions.
Figure 2 illustrates the SEC framework, integrating these components to learn personalized recommendation policies that optimize user retention. The following sections delve into the details of each component, describing the training and inference stages of the SEC framework.
Figure 3. T-SNE visualization of user state embeddings from the KuaiRand dataset, categorized by expert levels based on user return time. The plot shows the embedding space of level-1 experts, level-2 experts, level-3 experts, and non-expert users. Notably, some non-expert users align more closely with lower-level experts rather than the highest-retention group, motivating our multi-level stratification approach.
4.2. Training Stage
The training stage of the SEC framework involves multi-level expert stratification, behavior cloning for each expert level, action entropy regularization, and the optimization of training objectives.
4.2.1. Multi-Level Expert Stratification
The SEC framework introduces a multi-level expert stratification approach to capture the nuances in expert user behaviors at different retention levels. Our key insight is that even among expert users with high overall retention, there exist distinct subgroups exhibiting different behavioral patterns. As shown in Figure 3, some non-expert users do not necessarily fall within the representation space of the highest-retention experts, but rather align more closely with the second or third level of experts. This observation motivates our design of a hierarchical expert stratification scheme.
First, we define expert users as those who exhibit high engagement and long-term retention with the recommender system. The criterion for selecting expert users can be based on metrics such as the number of active days, sessions, or interactions within a given time period. Let $\mathcal{U}_{E}\subseteq\mathcal{U}$ denote the set of expert users.
Then we stratify the expert users into $K$ levels, $\{\mathcal{U}_{E}^{1},\mathcal{U}_{E}^{2},\ldots,\mathcal{U}_{E}^{K}\}$, based on their retention scores. The retention score can be computed using various metrics, such as the average number of active days per month or the lifetime value (LTV) of the user. The choice of the stratification metric and the number of levels $K$ can be determined based on domain knowledge and data analysis. For example, users can be divided into levels based on quantiles of the retention score distribution.
4.2.2. Behavior Cloning for Each Expert Level
For each expert level $k$, we learn a specialized behavior cloning policy $\pi_{k}$ that mimics the actions of experts in that level. The policy $\pi_{k}$ maps the user state $s_{t}$ to a recommendation action $a_{t}$, i.e., $\pi_{k}:\mathcal{S}\rightarrow\mathcal{A}$.
The behavior cloning objective for expert level $k$ is to minimize the negative log-likelihood of the expert actions given the corresponding user states:
(2) $\mathcal{L}_{\text{BC}}^{k}=\sum_{(s,a)\in\mathcal{D}_{k}}-\log\pi_{k}(a|s),$
where $\mathcal{D}_{k}=\{(s,a)\}$ is the dataset of state-action pairs collected from the interaction trajectories of expert users in level $k$.
We parameterize the policy $\pi_{k}$ using a neural network architecture that consists of two main components: a state encoder $f_{\phi}$ and an action predictor $g_{\theta_{k}}$. The state encoder $f_{\phi}$ maps the user state $s_{t}$ to a compact representation $h_{t}=f_{\phi}(s_{t})$, capturing the relevant features for decision making. The action predictor $g_{\theta_{k}}$ takes the state representation $h_{t}$ as input and generates a probability distribution over the action space $\mathcal{A}$.
For continuous action spaces, the action predictor can be modeled using a Gaussian distribution, and the behavior cloning objective can be formulated as minimizing the mean squared error (MSE) between the predicted and expert actions:
(3) $\mathcal{L}_{\text{BC}}^{k}=\sum_{(s,a)\in\mathcal{D}_{k}}\|g_{\theta_{k}}(f_{\phi}(s))-a\|^{2}.$
For discrete action spaces, the action predictor can be modeled using a categorical distribution, and the behavior cloning objective becomes the cross-entropy loss:
(4) $\mathcal{L}_{\text{BC}}^{k}=\sum_{(s,a)\in\mathcal{D}_{k}}-\log g_{\theta_{k}}(a|f_{\phi}(s)).$
The state encoder $f_{\phi}$ is shared across all expert levels, allowing for efficient learning and generalization. The action predictors $g_{\theta_{k}}$ are specific to each expert level, enabling specialized policies for different retention levels.
4.2.3. Action Entropy Regularization
Figure 4 illustrates the issue of limited diversity in the recommendation policy learned from behavior cloning. The t-SNE visualization of the action embeddings generated by the policy reveals a highly concentrated distribution. This lack of diversity can hinder the policy’s ability to capture users’ diverse interests and adapt to their evolving preferences. To encourage recommendation diversity and exploration, we introduce an Action Entropy Regularization (AER) term in the training objective. AER prevents the learned policies from collapsing to a narrow set of actions and promotes the discovery of novel and relevant recommendations.
In the case of continuous action spaces, such as when the recommendations are generated in the form of item embeddings, inspired by MMCR (Schaeffer et al., 2024), we propose to use the nuclear norm of the action matrix as the entropy. Let $A_{k}\in\mathbb{R}^{B\times d}$ denote the action matrix from $\pi_{k}$, where $B$ is batch size and $d$ is the embedding dimension. Then the regularization term is
$\mathcal{L}_{\text{AER}}^{k}=-|A_{k}|^{*}=-\sum_{i=1}^{\min(B,d)}\sigma_{i}(A_{k})$
where $\sigma_{i}(A_{k})$ denotes the $i$-th singular value of $A_{k}$.
In the case of discrete action spaces, where the recommendations are generated as categorical items, we compute the average probability for each class and maximize the entropy of the class probabilities:
$\mathcal{L}_{\text{AER}}^{k}=\sum_{j=1}^{C}p_{k,j}\log p_{k,j}$
where $p_{k,j}$ is the average probability of class $j$ for the $k$-th expert, and $C$ is the total number of classes.
4.2.4. Training Objectives
The overall training objective of the SEC framework is a combination of the behavior cloning loss and the action entropy regularization term, summed over all expert levels:
(5) $\mathcal{L}=\sum_{k=1}^{K}\left(\mathcal{L}_{\text{BC}}^{k}+\lambda\mathcal{L}_{\text{AER}}^{k}\right),$
where $\lambda$ is a hyperparameter that controls the strength of the regularization.
During training, the parameters of the state encoder $f_{\phi}$ and the action predictors $g_{\theta_{k}}$ are optimized using stochastic gradient descent (SGD) or its variants, such as Adam.
Figure 4. Visualization of the learned action distribution of a single expert policy in the principal component space (left) and t-SNE embedding space (right) using the base behavior cloning approach. The narrow concentration of actions in both projections reveals a lack of diversity in the recommended items. This limited coverage of the action space may hinder the policy’s ability to capture diverse user preferences, motivating the need for entropy regularization to encourage more diverse and exploratory recommendations.
4.3. Inference Stage
During the inference stage, the SEC framework leverages the trained expert-level policies and the adaptive selection mechanism to generate personalized recommendations for each user.
4.3.1. User State Representation and Clustering
To capture the representative user states for each expert level, we apply the K-means clustering algorithm to partition the compact state space from the state encoder $f_{\phi}$ of each expert level into $C_{k}$ clusters. The centroid of each cluster represents a typical user state for that expert level. Let $\mu_{k,c}$ denote the centroid of the $c$-th cluster in the $k$-th expert level.
4.3.2. Adaptive Expert Selection Mechanism
As shown in the right part of Figure 2, The adaptive expert selection mechanism dynamically assigns users to the most appropriate expert level based on their current state and historical interactions. This allows the recommender system to adapt to the evolving preferences and retention patterns of each user. Given a user state $s$, the adaptive selection mechanism computes the minimum distance between the user state and the cluster centroids of each expert level:
(6) $d_{k}=\min_{c}d(f_{\phi}(s),\mu_{k,c}),\quad k=1,2,\ldots,K$
The user is then assigned to the expert level $k^{*}$ that satisfies:
(7) $k^{*}=\text{argmin}_{k}\{k|d_{k}\leq\delta_{k}\}$
where $\delta_{k}$ is a predefined threshold for each expert level. If the user state does not match any expert level, the user is assigned to the expert level with the minimum distance. To incorporate the user’s historical retention information, we introduce a constraint on the selected expert level. Let $r_{h}$ denote the user’s historical retention level, estimated based on their past interactions and retention metrics. We restrict the selected expert level $k^{*}$ to be no lower than the historical retention level $r_{h}$:
(8) $k^{*}=\min(k^{*},r_{h})$
This constraint ensures that the selected expert level aligns with the user’s historical retention patterns and prevents the recommendation quality from degrading.
4.3.3. Generating Retention-Oriented Recommendations
During the inference stage, the SEC framework generates personalized recommendations optimized for user retention. Once the most suitable expert level $k^{*}$ is determined for a user based on their current state and historical interactions, the corresponding expert policy $\pi_{k^{*}}$ is applied to generate recommendations.
The expert policy $\pi_{k^{*}}$ takes the user state $s$ as input and outputs a set of recommended items or actions that are aligned with the retention strategies learned from the high-retention expert users at level $k^{*}$. By mimicking the actions of expert users who have demonstrated strong retention patterns, the policy aims to generate recommendations that are likely to keep the user engaged and promote long-term retention.
By adaptively selecting the most appropriate expert policy based on the user’s state and generating retention-oriented recommendations, the SEC framework tailors its recommendations to the specific preferences and retention patterns of each user. This personalized approach aims to improve long-term user retention and satisfaction by providing recommendations that are most likely to keep the user engaged and active on the platform.
Figure 5. Visualization of action distributions of one expert in the principal component space (left) and t-SNE embedding space (right) with action entropy regularization. The regularization promotes a more diverse set of actions compared to base behavior cloning.
5. Experiments
We evaluate the performance of our method SEC through extensive offline experiments and online A/B tests on two major video platforms with hundreds of millions of daily active users. During our offline testing, continuous actions are utilized, whereas discrete actions are used for the online experiments.
We aim to answer the following research questions through our experiments:
  • RQ1: How does SEC compare with state-of-the-art methods in offline evaluations?
  • RQ2: What are the effects of each component in SEC based on offline analysis?
  • RQ3: How does SEC perform in terms of user retention in online settings?
  • RQ4: Can SEC be extended to optimize other long-term values in online experiments?
5.1. Offline Experiments
Table 1. Overall Performance on KuaiRand-Pure dataset for different models.
Model Return Time Click Rate Long View Rate Like Rate
TD3 2.382 0.800 0.791 0.852
SAC 2.373 0.801 0.795 0.857
DIN 1.947 0.773 0.764 0.812
CEM 1.889 0.762 0.757 0.804
RLUR 1.786 0.789 0.778 0.831
GFN 1.496 0.805 0.794 0.885
SEC(Ours) 1.411 0.833 0.825 0.806
5.1.1. Experimental Setup
For offline evaluations, we utilize the KuaiRand dataset, a sequential recommendation dataset that provides an unbiased setting by incorporating random video exposures. KuaiRand encompasses a rich set of user feedback signals, including six positive signals (’click’, ’view time’, ’like’, ’comment’, ’follow’, and ’forward’) and two negative signals (’hate’ and ’leave’). To maintain focus and clarity in our analysis, we concentrate on the six positive signals and the two most relevant negative signals, while excluding less frequent feedback types. The KuaiRand dataset comprises interactions from 27,285 users on 7,551 items, totaling 1,436,609 interactions with a density of 0.70%.
To emulate realistic long-term user behavior in the context of short video recommendations, we employ the KuaiSim retention simulator (Zhao et al., 2023a). KuaiSim incorporates two primary modules: (1) a leave module that predicts the probability of a user exiting the current session, thus concluding an episode; and (2) a return module that estimates the daily likelihood of a user revisiting the platform, represented as a multinomial distribution. By leveraging KuaiSim, we can effectively assess the performance of recommendation algorithms in fostering long-term user engagement and retention. The dataset and KuaiSim retention configuration used in our experiments follow the settings described in the GFN model (Liu et al., 2024a).
In our offline experiments, We consider users with a return time less than or equal to 3 days as experts, and users with return times of one day, two days, and three days are considered experts at different levels. We employ the Policy network framework from GFN, which uses a Transformer architecture, for our behavior cloning approach. The user embeddings are pre-trained from the user model and fixed in our model. During inference, we set the number of clusters for each expert level to 1000 and define $\delta_{k}$ as half of the mean distance between clusters for each expert group. Optimization is carried out using the Adam optimizer with a batch size of 128. We apply a default balance parameter $\lambda$ of 0.01 and set the learning rate to 0.0003.
5.1.2. Baselines and Metrics
Our model is benchmarked against a diverse array of baselines, including:
  • CEM (Rubinstein and Kroese, 2004): A model-free optimization algorithm widely used in reinforcement learning tasks.
  • DIN (Zhou et al., 2018): A deep learning-based recommendation model that captures users’ dynamic interests from their historical behaviors.
  • TD3 (Fujimoto et al., 2018): An off-policy reinforcement learning algorithm that extends DDPG and addresses overestimation bias in Q-learning.
  • SAC (Haarnoja et al., 2018): An off-policy maximum entropy deep reinforcement learning algorithm that encourages exploration and robustness.
  • RLUR (Cai et al., 2023b): A reinforcement learning-based framework specifically designed for optimizing long-term user retention in recommender systems.
  • GFN (Liu et al., 2024a): A recently proposed generative model for sequential recommendation that learns a generative policy to model the probability distribution of user behavior sequences.
To evaluate the performance of the recommendation algorithms, we employ the following metrics:
  • Return Time: The average number of days it takes for a user to return to the platform after their last interaction. A lower return time indicates better user retention.
  • Click Rate: The ratio of the number of clicks to the total number of recommendations. A higher click rate suggests that the recommendations are more relevant to users’ interests.
  • Long View Rate: The proportion of video views that exceed a certain duration threshold (e.g., 50% of the video length). A higher long view rate implies that users find the recommended videos engaging and tend to watch them for a longer time.
  • Like Rate: The ratio of the number of likes to the total number of recommendations. A higher like rate indicates that users express more immediate satisfaction with the recommended items.
These metrics provide a comprehensive assessment of the recommendation algorithms’ performance in terms of user retention (Return Time), user engagement (Click Rate and Long View Rate), and user satisfaction (Like Rate). By evaluating the models across these diverse metrics, we can gain insights into their effectiveness in optimizing long-term user retention and delivering a satisfactory user experience.
Table 2. Ablation Study Results on KuaiRand-Pure dataset.
Variant Return Day Click Rate Long View Rate
SEC (Full) 1.411 0.833 0.825
w/o Multi-level 1.423 0.821 0.815
w/o Action Entropy 1.449 0.836 0.805
5.1.3. RQ1 Results
Figure 6. The performance of SEC with different balance loss factors $\gamma$ on the KuaiRand Dataset for minimizing return time.
Table 1 presents the comprehensive offline experimental results, comparing the performance of SEC against various baseline methods. SEC demonstrates superior performance across most metrics, particularly in terms of Return Time, Click Rate, and Long View Rate. Focusing on the Return Time metric, which directly reflects user retention, SEC achieves a significant reduction of 5.7% compared to the best baseline model, GFN. This improvement indicates that SEC effectively prolongs user lifecycles and encourages users to return to the platform more frequently. Similarly, SEC outperforms GFN in terms of Click Rate and Long View Rate, with improvements of 3.5% and 3.9%, respectively. These results suggest that SEC not only enhances user retention but also increases user engagement, as evidenced by higher click-through rates and longer view durations.
However, it is worth noting that SEC’s performance in terms of Like Rate is lower compared to GFN, with a difference of 8.9%. This can be attributed to the primary objective of SEC, which is to optimize long-term user retention. In pursuit of this goal, SEC may prioritize recommendations that encourage users to stay on the platform and explore more content, rather than solely focusing on maximizing immediate user satisfaction, as indicated by the Like Rate.
It is important to recognize that user retention and immediate user satisfaction are distinct objectives, and there may be trade-offs between them. While a high Like Rate is generally desirable, it does not necessarily translate to long-term user retention. Users may express immediate satisfaction with a recommendation by liking it, but this does not guarantee their continued engagement with the platform over an extended period.
5.1.4. RQ2 Results
We conduct an ablation study to analyze the contribution of each component in SEC. Table 2 presents the results. Removing the multi-level behavior cloning leads to a 0.9% increase in Return Time and a over 1.0% decrease in both Click Rate and Long View Rate, indicating the importance of learning from experts at different retention levels in improving user retention and engagement. When removing the action entropy regularization, we observe a 1.8% increase in Return Time, similar with the overall goal of SEC to optimize long-term user retention. However, the removal of action entropy regularization also results in a 1.7% increase in Click Rate, accompanied by a 2.4% decrease in Long View Rate, suggesting that while the absence of action entropy regularization may lead to higher immediate user engagement, it may not necessarily translate to sustained engagement.
To further investigate the impact of the balance factor between behavior cloning and action entropy regularization, we conduct experiments with different values of the balance factor and evaluate their effect on the return time. Figure 6 presents the results, revealing an interesting pattern. When the balance factor is set to a low value, indicating a lower emphasis on action entropy regularization, the return time is relatively high, around 1.45. As the balance factor increases, the return time initially improves, reaching its optimal value of 1.411 at a moderate balance factor setting. However, when the balance factor becomes too large, prioritizing action entropy regularization over behavior cloning, the performance deteriorates significantly, with the return time increasing to 1.58. This observation highlights the importance of striking the right balance between behavior cloning and action entropy regularization in SEC.
To further investigate the impact of action entropy regularization, we visualize the learned action representations using t-SNE, as shown in Figure 5. The visualization reveals that the action embeddings learned by SEC exhibit significant diversity, with a high degree of entropy. Rather than forming distinct clusters, the actions are spread out in the embedding space, indicating that SEC learns a wide range of recommendation strategies.
5.2. Online Experiments
Figure 7. Online experimental environment, consistent with GFN (Liu et al., 2024a), where the SEC model collaborates with other strategies to optimize user retention.
5.2.1. Experimental Setup
In the online experiments, we deploy SEC on two major video platforms, Kuaishou and Kuaishou Lite, with hundreds of millions of daily active users. In the offline experiments, our strategy has full control over the ranking of items, allowing the actions to be combinations of multiple predictive factors, resulting in a continuous action space. Nevertheless, as demonstrated in Figure 7, during online experiments, we must incorporate the recommendation rankings of other scores that are beyond our control. Given the large scale of video IDs, directly using them as actions would be computationally infeasible. Therefore, we employed video clustering. Specifically, we clustered videos based on multimodal features, grouping similar videos together. The recommendation actions are then generated based on these clusters, which not only improves computational efficiency but also enhances the diversity and relevance of the recommendations. State Encoder and Action Predictor both use MLP network. We used users with high activity days as experts and classify them into 3 levels. During online recommendation, we combine cluster action prediction probability with other fine-grained ranking scores, which already include GFN (Liu et al., 2024a) as a baseline, and apply them in the ranking stages, as shown in Figure 7.
We conduct online A/B tests, optimizing for long-term user retention. In the tests, we allocate 10% of the real traffic to the baseline version and the optimized version (using SEC) and run the experiments for over two weeks.
5.2.2. RQ3 Results
The experiments focus on user retention, and the results are presented in Table 3. We use the LT7 metric to calculate Active Days, defined as the number of active users within a 7-day window divided by the total number of users in the group.
Table 3. Online A/B test results for User Retention. Bold values indicate statistically significant results ($p$-value $<5\%$).
Model Stage Kuaishou Kuaishou Lite
Behavior Cloning + Entropy Loss 1st Ranking +0.034% +0.037%
2nd Ranking +0.036% +0.042%
Additional gain from Multi-level Expert Both Stages +0.028% +0.043%
Total Improvement +0.098% +0.122%
The experimental results demonstrate that SEC achieves significant improvements in Active Days across both platforms. Initially, we implemented a model with Behavior Cloning + Entropy Loss, which showed promising results. For the 1st Ranking stage, Kuaishou showed a gain of +0.034%, while Kuaishou Lite exhibited an improvement of +0.037%. In the 2nd Ranking stage, these gains increased to +0.036% for Kuaishou and +0.042% for Kuaishou Lite.
After observing these positive results, we enhanced our approach by incorporating the Multi-level Expert stratification component of SEC, replacing the previous model in both ranking stages. This further improvement led to additional gains of +0.028% on Kuaishou and +0.043% on Kuaishou Lite.
Overall, the cumulative improvements are +0.098% for Kuaishou and +0.122% for Kuaishou Lite, additionally bringing tens of thousands of daily active users to each platform. These results validate the effectiveness of the complete SEC framework in prolonging user lifecycles and enhancing user retention.
Additionally, we examine the impact of incorporating action entropy regularization into our training process. The results indicate that this addition improves AUC scores during online training, with varying levels of improvement observed across different activity levels of experts, as shown in Figure LABEL:fig:auc_results. Notably, our AUC calculations are based on a dataset containing billions of samples. This suggests that our action entropy regularization and behavior cloning loss are isotropic, effectively enhancing overall model performance.
Figure 8. AUC scores during online experiments with and without action entropy regularization across different expert activity levels.
5.2.3. RQ4 Results
To further demonstrate the effectiveness of our approach, we conduct an additional set of experiments focusing on user interest expansion. The results are presented in Table 4.
Table 4. Online A/B test results for User Interest Expansion. Bold values indicate statistically significant results ($p$-value $<5\%$).
Metric Stage Kuaishou Kuaishou Lite
Valid Clusters 1st Ranking +1.31% +1.14%
We evaluate the impact of SEC on the number of valid interest clusters (Valid Clusters) for users. Valid interest clusters refer to the video clusters that users frequently interact with within a certain time period. This metric indirectly reflects the breadth of user interactions and the diversity of their interests on the platform. The experimental results show that SEC increases the Valid Clusters metric by 1.31% on Kuaishou and 1.14% on Kuaishou Lite.
The online experiments highlight the practical value of SEC in industrial-scale recommendation systems. By learning from the behaviors of experts at different retention levels, adaptively selecting the best expert strategies, and encouraging diversity in recommendation actions, SEC achieves notable improvements in both user retention and user interest expansion, depending on the optimized long-term value. The use of clustering to handle the vast scale of video content not only reduces computational complexity but also enhances the generalizability and robustness of the recommendation system, providing valuable insights for engineering practices in large-scale recommendations.
6. Conclusion
We propose Stratified Expert Cloning (SEC), a novel imitation learning framework for optimizing long-term user retention in large-scale recommender systems. SEC leverages multi-level behavior cloning to learn specialized policies for expert users at different retention levels, an adaptive expert selection mechanism to assign users to the most appropriate level, and action entropy regularization to promote recommendation diversity and exploration.
Extensive offline experiments demonstrate SEC’s superiority over state-of-the-art methods in improving retention metrics. Our online A/B tests on two major video platforms, Kuaishou and Kuaishou Lite, validate the practical impact of our approach with cumulative lifts of 0.098% and 0.122% in active days respectively, additionally bringing tens of thousands of daily active users to each platform. Our ablation studies confirm the effectiveness of each SEC component, with the multi-level expert stratification providing significant additional gains beyond basic behavior cloning with entropy regularization.
This work highlights the promise of imitation learning for retention-aware recommender systems in production environments. Future directions include incorporating inverse reinforcement learning, semi-supervised learning, and transfer learning. SEC lays a foundation for developing recommenders that deliver long-lasting user value at scale, with proven business impact across multiple platforms.